• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:02:19 on July 2, 2001

    Entry number 65455

    keyword=Beam Time Accounting
    Hall A

    7/2/2001 0800-1600
    Experiment #: e99117
    Shift Leader: A. Deur
    Shift Workers: S. Binet, J. Singh

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    34 30 49 45 38 46 53 53 5.8
    Beam Available
    but Not in Use
    12 0 0 0 0 0 0 0 0.2
    Beam Not Available
    or Unacceptable
    14 30 11 15 22 14 7 7 2
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 1 1 1 1 1 1 8
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 12 0 0 0 0 0 0 0 0.2
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Ac 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 8:38. Ask MCC to turn off the beam in order to perform a target polarimetry (NMR). 8:50. Ask for the beam back 8:55 Beam is still down due to cavity problem. 9:30 Beam is back. We resume data taking. 12:55 Beam is down 13:10. Beam is back. 2:10. EPR measurement. 2:25. Ask for the beam back at 12 uA.

    Data Entry by: A. Deur